OpenStudy (anonymous):

Verify the identity. Show your work. 1 + sec^2xsin^2x = sec^2x my work: os^2x+sin^2x=1 so (if you divide both sides by cos^2x)... 1+tan^2x=sec^2x tan^2x=sec^2x-1 sec^2x/sec^2x-1=csc^2x sec^2x/tan^2x=csc^2x (1/cos^2x)/(sin^2x/cos^2x)=csc^2x (1/cos^2x)(cos^2x/sin^2x)=csc^2x (1/sin^2x)=csc^2x csc^2x=csc^2x it has some error what i dont know @mathmath333

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ \color{red}{ 1 + sec^2xsin^2x = sec^2x} \\~\\ 1 + sec^2xsin^2x \\~\\ =1 + \dfrac{1}{cos^2x}sin^2x \\~\\ =\dfrac{cos^2x + sin^2x}{cos^2x} \\~\\ =\dfrac{1}{cos^2x} \\~\\ =sec^2x }\end{align}\)

OpenStudy (anonymous):

3) Find the vertical asymptotes 2, if any, of the graph of the rational function. Show your work. f(x) = (x-4)/(x(x-4)) Vertical asymptotes: x-4 F(x) = ______ X(x-4) x-4 F(x) = ______  taking common out X(x-4) Leaves it f(x)=x So the answer would be Vertical asymptotes = x=0, or x=4 find the VA's by setting the denominator to 0. When we get a factor that cancels out, it is not a vertical asymptote. What is it? Graph it to see what it looks like.

OpenStudy (jhannybean):

\[1 + \sec^2x\sin^2x = \sec^2x \]\[1+\frac{\sin^2(x)}{\cos^2(x)}=\sec^2(x)\]\[1+\tan^2(x)=\sec^2(x)\]\[\sec^2(x)=\sec^2(x)~\checkmark\]

OpenStudy (jhannybean):

Known identities:\[\sin^2(x) +\cos^2(x) = 1\]\[1+\tan^2(x) = \sec^2(x)\]\[1+\cot^2(x) = \csc^2(x)\]

OpenStudy (jhannybean):

For 3. Vertical asymptotes happen when denominator = 0

Parth (parthkohli):

\[1 + \sec^2 x (1 - \cos^2 x) = 1 + \sec^2 x - 1 = \sec^2 x\]So many ways...

OpenStudy (jhannybean):

Mmhmm, trig is wonderful.

OpenStudy (anonymous):

thanx dear @mathmath333 already help me in that one can u help me solve other Q's?

OpenStudy (anonymous):

@Jhannybean i am following ur explanation on Q-3

OpenStudy (mathmath333):

\[x(x-4)=0\\\text{solve forvertical asymptote}\]

OpenStudy (mathmath333):

that is the denominator part

OpenStudy (jhannybean):

Sorry, distracted.

OpenStudy (mathmath333):

there should be 2 asymptote

OpenStudy (anonymous):

yes i solve it and got x=0 and x=4

OpenStudy (jhannybean):

Yeah, that's what I got as well.

OpenStudy (anonymous):

but my teacher said : You find the VA's by setting the denominator to 0. When we get a factor that cancels out, it is not a vertical asymptote. What is it? Graph it to see what it looks like. what i did wrong?

OpenStudy (anonymous):

@jhannybean @mathmath333

OpenStudy (anonymous):

@Nnesha could u help me in above question

OpenStudy (jhannybean):

a factor that cancels out is what you call a removable discontinuity,

OpenStudy (anonymous):

so what step is missing plzz help